# A lower bound for the Weisfeiler-Leman dimension of circulant graphs

### v1

@@ 7. Local system of multipliers
To prove the second equality in (33), we need to verify that ${\mathscr{X}}^{f}\in\operatorname{WL}_{\ell}({\cal X})$ for each ${\mathscr{X}}\in\operatorname{WL}_{\ell}({\cal X})$. Note that by virtue of (30), there is a class ${\mathscr{X}}^{\prime}\in{\cal R}_{\ell}({\cal X})$ containing the class ${\mathscr{X}}$. Therefore, ${\mathscr{X}}^{f}={\mathscr{X}}^{f_{{\mathscr{X}}^{\prime}}}$. Since $f_{{\mathscr{X}}^{\prime}}\in\operatorname{Iso}({\cal X})$, we conclude that ${\mathscr{X}}^{f_{{\mathscr{X}}^{\prime}}}\in\operatorname{WL}_{\ell}({\cal X})$. Thus, ${\mathscr{X}}^{f}\in\operatorname{WL}_{\ell}({\cal X})$, as required. ∎
@@
Denote by $\hat{f}:G^{m+1}\to G^{m+1}$ the bijection $f$ defined in Lemma 7.2 for $\ell=m+1$; note that by the assumption on $m$, we have $\ell\leq\lfloor\sqrt{k}\rfloor$. It suffices to prove that if $\hat{\mathscr{X}}\in{\cal R}_{m+1}({\cal X})$ and $x\in\operatorname{pr}_{m}\hat{\mathscr{X}}$, then there exists a bijection $\theta_{\hat{\mathscr{X}}}:\operatorname{Res}_{x}\hat{\mathscr{X}}\to\operatorname{Res}_{x^{f}}\hat{\mathscr{X}}^{\hat{f}}$ such that
@@
Indeed, $G$ is a disjoint union of those sets $\operatorname{Res}_{x}\hat{\mathscr{X}}$ for which $\hat{\mathscr{X}}\in{\cal R}_{m+1}({\cal X})$ and $x\in\operatorname{pr}_{m}\hat{\mathscr{X}}$. Moreover, assuming the existence of the bijections $\theta_{\hat{\mathscr{X}}}$, we see that
@@
To prove formula (35), fix $\alpha\in\operatorname{Res}_{x}\hat{\mathscr{X}}$ and put ${\mathscr{X}}=\operatorname{pr}_{m}\hat{\mathscr{X}}$. Let $\hat{\mathfrak{S}}={\mathfrak{S}}(\hat{\mathscr{X}})$ and ${\mathfrak{S}}={\mathfrak{S}}({\mathscr{X}})$ be the sets of principal ${\cal A}$-sections, defined by formula (32). Then ${\mathfrak{S}}\subseteq\hat{\mathfrak{S}}$, and ${\mathfrak{S}},\hat{\mathfrak{S}}\in{\mathfrak{S}}_{k}$ by the choice of $m$. Since $x\in{\mathscr{X}}$ and $x\cdot\alpha\in\hat{\mathscr{X}}$, this implies that
@@
According to our definition, $f_{\mathscr{X}}=f_{{\mathfrak{m}}({\mathscr{X}})}$ and $\hat{f}_{\hat{\mathscr{X}}}=f_{{\mathfrak{m}}(\hat{\mathscr{X}})}$ are the isomorphisms of the S-ring ${\cal A}$, inducing the inner ${\cal A}$-multipliers ${\mathfrak{m}}({\mathfrak{S}})$ and ${\mathfrak{m}}(\hat{\mathfrak{S}})$, respectively. Now the left-hand side of formula (28) implies that
@@
for all principal ${\cal A}$-sections ${\mathfrak{s}}_{ij}={\mathfrak{s}}_{X_{ij}}$ belonging to the subset $\hat{\mathfrak{S}}$. This implies that $(X_{ij})^{\hat{f}}=\varphi(X_{ij})$. By the right-hand side of formula (36), this yields
@@
At this point, we need an auxiliary statement concerning the array $R_{0}(\cdot)=R_{{\cal X}_{0},m}(\cdot)$ defined by formula (7), where ${\cal X}_{0}={\cal X}({\cal A}_{0})$ (recall that ${\cal A}_{0}$ is the coset closure of ${\cal A}$).
@@
Let $h=h({\mathfrak{S}},\hat{\mathfrak{S}})$ be the automorphism of the S-ring ${\cal A}$, corresponding to the inner ${\cal A}$-multiplier ${\mathfrak{m}}({\mathfrak{S}},\hat{\mathfrak{S}})\in{\mathfrak{M}}_{aut}({\cal A})$ on the right-hand side of formula (28). Then
@@
We need to verify that for all $1\leq i,j\leq m$, the relations $r_{0}((x^{\hat{f}})_{i},(x^{\hat{f}})_{j})$ and $r_{0}((x^{fh})_{i},(x^{fh})_{j})$ are equal, where $r_{0}=r_{{\cal X}_{0}}$. Recall that $\hat{f}_{\hat{\mathscr{X}}}\in\operatorname{Iso}_{\cal A}({\cal A}_{0})$ and hence $\hat{f}_{\hat{\mathscr{X}}}\in\operatorname{Iso}({\cal X}_{0})$. Therefore,
@@
Similarly, since $f_{\mathscr{X}}$ and $h_{\mathscr{X}}$ are also isomorphisms of the scheme ${\cal X}_{0}$, we have
@@
It remains to verify that the relations on the right-hand sides of formulas (38) and (39) coincide. Denote by $X_{0}$ the basic set of the S-ring ${\cal A}_{0}$ that corresponds to the basis relation $r_{0}(x_{i},x_{j})$. Then $X_{0}$ contains the element $x_{0}=x_{j}x_{i}^{-1}$. The element $\bar{x}_{0}=(x_{0})_{\mathfrak{s}}$ belongs to the set $X_{ij}$ that generates the cyclic group ${\mathfrak{s}}={\mathfrak{s}}_{X_{ij}}$. Since the S-ring ${\cal A}$ is totally normal, the S-ring ${\cal A}_{\mathfrak{s}}$ is normal and hence every element of $X_{ij}$ is a generator of ${\mathfrak{s}}$; in particular, ${\mathfrak{s}}=\langle{x_{0}}\rangle$. It follows that ${\mathfrak{s}}={\mathfrak{s}}_{X_{0}}$. Consequently,
@@
where $\hat{\mathfrak{m}}={\mathfrak{m}}(\hat{\mathfrak{S}})$, ${\mathfrak{m}}={\mathfrak{m}}({\mathfrak{S}})$, ${\mathfrak{n}}={\mathfrak{m}}({\mathfrak{S}},\hat{\mathfrak{S}})$. and the equality $\hat{\mathfrak{m}}_{\mathfrak{s}}=m_{\mathfrak{s}}n_{\mathfrak{s}}$ follows from the second equality in (28). Therefore,
@@
By theorem condition, the scheme ${\cal X}_{0}$ is binary. Besides, $R_{0}(x^{\hat{f}})=R_{0}(x^{fh})$ by Claim 1, where $h\in\operatorname{Aut}({\cal A})$. Thus there exists $h_{0}\in\operatorname{Aut}({\cal X}_{0})\leq\operatorname{Aut}({\cal X})$ such that $x^{\hat{f}h_{0}}=x^{fh}$. Then
@@
Note that $x^{\hat{f}}\cdot\alpha^{\hat{f}}\in{\hat{\mathscr{X}}}^{\widehat{f}}$ by the choice of $\alpha$. Since $h_{0}$ and $h^{-1}$ are automorphisms of the scheme ${\cal X}$, this implies that $x^{f}\cdot\alpha^{\theta_{\hat{\mathscr{X}}}}\in{\hat{\mathscr{X}}}^{\widehat{f}}$, where $\theta_{\hat{\mathscr{X}}}=\hat{f}h_{0}h^{-1}$ is the composition of isomorphisms of ${\cal X}$. In particular,
@@
To complete the proof, we recall that the point $\alpha$ was chosen in $\operatorname{Res}_{x}\hat{\mathscr{X}}$ arbitrarily. Moreover, none of the isomorphisms $\hat{f}$, $h_{0}$, and $h$ depends on $\alpha$, and the mapping $\theta_{\hat{\mathscr{X}}}:\operatorname{Res}_{x}\hat{\mathscr{X}}\to\operatorname{Res}_{x^{f}}\hat{\mathscr{X}}^{\hat{f}}$ is a bijection. Finally, by formulas (37) and (40) and because of $h_{0},h\in\operatorname{Aut}({\cal X})$, we have
$R(x\cdot\alpha)^{\varphi}=R(x^{\hat{f}}\cdot\alpha^{\hat{f}})=R(x^{\hat{f}}\cdot\alpha^{\hat{f}})^{h_{0}h^{-1}}=R(x^{f}\cdot\alpha^{\hat{f}h_{0}h^{-1}})=R(x^{f}\cdot\alpha^{\theta_{\hat{\mathfrak{X}}}})$
@@
By Lemma 4.1, to prove that the scheme ${\cal X}$ is $\operatorname{WL}_{m}$-equivalent to itself with respect to the algebraic isomorphism $\varphi$, it suffices to verify that the bijection $f$ defined in Lemma 7.2 for $\ell=m$ satisfies the condition of Lemma 4.1. In other words, we need to verify that if $x\in G^{m}$ and $x^{\prime}=x^{f}$, then Duplicator has a winning strategy in the pebble game ${\mathscr{C}}_{m+1}(\varphi)$ on the scheme ${\cal X}$ with initial configuration $(x,x^{\prime})$. Let Spoiler choose a set $B\subseteq G$. Duplicator responds with the set $A=B^{\theta}$, where $\theta$ is the bijection defined in Lemma 7.3. Spoiler places one of the pebbles on a point $\alpha\in A$. Duplicator places the corresponding pebble to the point $b=\alpha^{\theta}$ and wins, because $R(x\cdot\alpha)^{\varphi}=R(x^{\prime}\cdot b)$ by formula (34). ∎
@@ 8. Two ingredients for the main construction
In the sequel, $a=2b$ is a positive integer and $V$ is a set of cardinality $a$. We fix a simple graph ${\cal G}=(V,E)$ belonging to the class ${\mathfrak{B}}_{a}$ defined in Section 1; in particular, $|E|=3a/2$. It is assumed that $V=\{v_{1},\ldots,v_{a}\}$. The vertices of the parts $V_{1}$ and $V_{2}$ of the (bipartite) graph ${\cal G}$ are denoted by $u_{1}:=v_{1},\ldots,u_{b}:=v_{b}$ and $w_{1}:=v_{b+1},\ldots,w_{b}:=v_{a}$, respectively.
The graph ${\cal G}$ being cubic and bipartite admits a $1$-factorization $F_{\cal G}$, i.e., the set of three pairwise disjoint matchings the union of which is equal to $E$. It is convenient to index the matchings of $F_{\cal G}$ by the nonidentity elements of the Klein group $K=C_{2}\times C_{2}$; we set $F_{\cal G}=\{E_{k}:\ k\in K^{\#}\}$, where $K^{\#}=K\setminus\{1\}$ and $E_{k}$ is the matching indexed by $k$.
8.1. A coherent configuration associated with ${\cal G}$
With each vertex $v\in V$ of the graph ${\cal G}$, we associate a copy $\Omega_{v}$ of the Klein group $K$ and, for a subset $V^{\prime}\subseteq V$, we denote by $\Omega_{V^{\prime}}$ the union of $\Omega_{v}$, $v\in V^{\prime}$. When $V^{\prime}=\{v,v^{\prime}\}$, we abbreviate $\Omega_{v,v^{\prime}}:=\Omega_{\{v,v^{\prime}\}}$.
@@ 8.1. A coherent configuration associated with ${\cal G}$
is a permutation group on the set $\Omega=\Omega_{V}$. It was proved in [8, Section 5] that there exists a coherent configuration ${\cal X}={\cal X}({\cal G},F_{\cal G})=(\Omega,S)$ that satisfies the following conditions:
@@
    $F({\cal X})=\{\Omega_{v}:\ v\in V\}$, and also $\operatorname{Aut}({\cal X})^{\Omega_{v}}=\bar{K}_{v}$ for all $v$,
@@
    the group of all $f\in\operatorname{Iso}({\cal X})$ such that ${\Omega_{v}}^{f}=\Omega_{v}$ and $f^{\Omega_{v}}\in\bar{K}_{v}$ for all $v\in V$, coincides with $K_{V}$,
@@
    $\operatorname{Aut}({\cal X})^{\Omega_{v,v^{\prime}}}=\bar{K}_{v}\times\bar{K}_{v^{\prime}}$ for all distinct $v,v^{\prime}\in V$ nonadjacent in ${\cal G}$,
@@
    $\operatorname{Aut}({\cal X})^{\Omega_{v,v^{\prime}}}=\{(x,x^{\prime})\in\bar{K}_{v}\times\bar{K}_{v^{\prime}}:\ \bar{\rho}_{v}(x)=\bar{\rho}_{v^{\prime}}(x^{\prime})\},$
@@
    if $\{u,w\}\in E$ and $\psi\in\operatorname{Sym}(S)$ are such that $\psi(s)\neq s$ if and only if $s\in S_{\Omega_{u},\Omega_{w}}\,\cup\,S_{\Omega_{w},\Omega_{u}}$, then $\psi$ is an algebraic automorphism of ${\cal X}$.
The details concerning the construction and other properties of the coherent configuration ${\cal X}$ can also be found in [5, Subsections 4.1.3 and 4.2.1], where ${\cal X}$ is defined as a special case of a cubic Klein configuration.
To state the main fact about the coherent configuration ${\cal X}$, that we need later, we fix an edge $e_{0}\in E$ and set $\psi=\psi_{e_{0}}$ to be the algebraic isomorphism in condition (K5) for $e=e_{0}$. Next, for an integer $m\geq 1$, we put
$F_{m}=F_{m}({\cal X})=\{\Omega_{V^{\prime}}:\ |V^{\prime}|\leq m\}.$
A family $\{f_{\Delta}\in K_{V}:\Delta\in F_{m}\}$ is called an $m$-local system of isomorphisms for ${\cal X}$ with respect to $\psi$ if for all $\Delta,\Delta^{\prime}\in F_{m}$, we have
$(f_{\Delta})^{\Delta}\in\operatorname{Iso}({\cal X}_{\Delta},\psi_{\Delta})\quad\text{and}\quad(f_{\Delta}{f_{\Delta^{\prime}}}^{-1})^{\Delta\cap\Delta^{\prime}}\in\operatorname{Aut}({\cal X})^{\Delta\cap\Delta^{\prime}}.$
@@
In the above notation, assume that each separator [^5] A separator of the graph ${\cal G}$ is a subset $V^{\prime}\subset V$ satisfying the following condition: each connected component of the graph obtained from ${\cal G}$ by removing all the vertices of $V^{\prime}$ contains at most half vertices of ${\cal G}$. of the graph ${\cal G}$ has cardinality greater than $3m$. Then $\operatorname{Iso}({\cal X},\psi)=\varnothing$ and there is an $m$-local system of isomorphisms for ${\cal X}$ with respect to $\psi$.
@@
8.2. A coset S-ring associated with the graph ${\cal G}$

### v2

@@ 7. Local system of multipliers
To prove the second equality in (33), we need to verify that ${\mathscr{X}}^{f}\in\operatorname{WL}_{\ell}({\mathcal{X}})$ for each ${\mathscr{X}}\in\operatorname{WL}_{\ell}({\mathcal{X}})$. Note that by virtue of (30), there is a class ${\mathscr{X}}^{\prime}\in{\mathcal{R}}_{\ell}({\mathcal{X}})$ containing the class $\mathscr{X}$}. Therefore, ${\mathscr{X}}^{f}={\mathscr{X}}^{f_{{\mathscr{X}}^{\prime}}}$. Since $f_{{\mathscr{X}}^{\prime}}\in\operatorname{Iso}({\mathcal{X}})$, we conclude that ${\mathscr{X}}^{f_{{\mathscr{X}}^{\prime}}}\in\operatorname{WL}_{\ell}({\mathcal{X}})$. Thus, ${\mathscr{X}}^{f}\in\operatorname{WL}_{\ell}({\mathcal{X}})$, as required. ∎
@@
Denote by $\hat{f}:G^{m+1}\to G^{m+1}$ the bijection $f$ defined in Lemma 7.2 for $\ell=m+1$; note that by the assumption on $m$, we have $\ell\leq\lfloor\sqrt{k}\rfloor$. It suffices to prove that if $\hat{\mathscr{X}}\in{\mathcal{R}}_{m+1}({\mathcal{X}})$ and $x\in\operatorname{pr}_{m}\hat{\mathscr{X}}$, then there exists a bijection $\theta_{\hat{\mathscr{X}}}:\operatorname{Res}_{x}\hat{\mathscr{X}}\to\operatorname{Res}_{x^{f}}\hat{\mathscr{X}}^{\hat{f}}$ such that
@@
Indeed, $G$ is a disjoint union of those sets $\operatorname{Res}_{x}\hat{\mathscr{X}}$ for which $\hat{\mathscr{X}}\in{\mathcal{R}}_{m+1}({\mathcal{X}})$ and $x\in\operatorname{pr}_{m}\hat{\mathscr{X}}$. Moreover, assuming the existence of the bijections $\theta_{\hat{\mathscr{X}}}$, we see that
@@
[…]
@@
[…]
@@
for all principal $\mathcal{A}$}-sections ${\mathfrak{s}}_{ij}={\mathfrak{s}}_{X_{ij}}$ belonging to the subset $\hat{\mathfrak{S}}$. This implies that $(X_{ij})^{\hat{f}}=\varphi(X_{ij})$. By the right-hand side of formula (36), this yields
@@
At this point, we need an auxiliary statement concerning the array $R_{0}(\cdot)=R_{{\mathcal{X}}_{0},m}(\cdot)$ defined by formula (7), where ${\mathcal{X}}_{0}={\mathcal{X}}({\mathcal{A}}_{0})$ (recall that ${\mathcal{A}}_{0}$ is the coset closure of $\mathcal{A}$}).
@@
[…]
@@
We need to verify that for all $1\leq i,j\leq m$, the relations $r_{0}((x^{\hat{f}})_{i},(x^{\hat{f}})_{j})$ and $r_{0}((x^{fh})_{i},(x^{fh})_{j})$ are equal, where $r_{0}=r_{{\mathcal{X}}_{0}}$. Recall that $\hat{f}_{\hat{\mathscr{X}}}\in\operatorname{Iso}_{\mathcal{A}}({\mathcal{A}}_{0})$ and hence $\hat{f}_{\hat{\mathscr{X}}}\in\operatorname{Iso}({\mathcal{X}}_{0})$. Therefore,
@@
Similarly, since $f_{\mathscr{X}}$ and $h_{\mathscr{X}}$ are also isomorphisms of the scheme ${\mathcal{X}}_{0}$, we have
@@
It remains to verify that the relations on the right-hand sides of formulas (38) and (39) coincide. Denote by $X_{0}$ the basic set of the S-ring ${\mathcal{A}}_{0}$ that corresponds to the basis relation $r_{0}(x_{i},x_{j})$. Then $X_{0}$ contains the element $x_{0}=x_{j}x_{i}^{-1}$. The element $\bar{x}_{0}=(x_{0})_{\mathfrak{s}}$ belongs to the set $X_{ij}$ that generates the cyclic group ${\mathfrak{s}}={\mathfrak{s}}_{X_{ij}}$. Since the S-ring $\mathcal{A}$} is totally normal, the S-ring ${\mathcal{A}}_{\mathfrak{s}}$ is normal and hence every element of $X_{ij}$ is a generator of $\mathfrak{s}$}; in particular, ${\mathfrak{s}}=\langle{x_{0}}\rangle$. It follows that ${\mathfrak{s}}={\mathfrak{s}}_{X_{0}}$. Consequently,
@@
[…]
@@
By theorem condition, the scheme ${\mathcal{X}}_{0}$ is binary. Besides, $R_{0}(x^{\hat{f}})=R_{0}(x^{fh})$ by Claim 1, where $h\in\operatorname{Aut}({\mathcal{A}})$. Thus there exists $h_{0}\in\operatorname{Aut}({\mathcal{X}}_{0})\leq\operatorname{Aut}({\mathcal{X}})$ such that $x^{\hat{f}h_{0}}=x^{fh}$. Then
@@
Note that $x^{\hat{f}}\cdot\alpha^{\hat{f}}\in{\hat{\mathscr{X}}}^{\widehat{f}}$ by the choice of $\alpha$. Since $h_{0}$ and $h^{-1}$ are automorphisms of the scheme $\mathcal{X}$}, this implies that $x^{f}\cdot\alpha^{\theta_{\hat{\mathscr{X}}}}\in{\hat{\mathscr{X}}}^{\widehat{f}}$, where $\theta_{\hat{\mathscr{X}}}=\hat{f}h_{0}h^{-1}$ is the composition of isomorphisms of $\mathcal{X}$}. In particular,
@@
To complete the proof, we recall that the point $\alpha$ was chosen in $\operatorname{Res}_{x}\hat{\mathscr{X}}$ arbitrarily. Moreover, none of the isomorphisms $\hat{f}$, $h_{0}$, and $h$ depends on $\alpha$, and the mapping $\theta_{\hat{\mathscr{X}}}:\operatorname{Res}_{x}\hat{\mathscr{X}}\to\operatorname{Res}_{x^{f}}\hat{\mathscr{X}}^{\hat{f}}$ is a bijection. Finally, by formulas (37) and (40) and because of $h_{0},h\in\operatorname{Aut}({\mathcal{X}})$, we have
[…]
@@
By Lemma 4.1, to prove that the scheme $\mathcal{X}$} is $\operatorname{WL}_{m}$-equivalent to itself with respect to the algebraic isomorphism $\varphi$, it suffices to verify that the bijection $f$ defined in Lemma 7.2 for $\ell=m$ satisfies the condition of Lemma 4.1. In other words, we need to verify that if $x\in G^{m}$ and $x^{\prime}=x^{f}$, then Duplicator has a winning strategy in the pebble game ${\mathscr{C}}_{m+1}(\varphi)$ on the scheme $\mathcal{X}$} with initial configuration $(x,x^{\prime})$. Let Spoiler choose a set $B\subseteq G$. Duplicator responds with the set $A=B^{\theta}$, where $\theta$ is the bijection defined in Lemma 7.3. Spoiler places one of the pebbles on a point $\alpha\in A$. Duplicator places the corresponding pebble to the point $b=\alpha^{\theta}$ and wins, because $R(x\cdot\alpha)^{\varphi}=R(x^{\prime}\cdot b)$ by formula (34). ∎
@@ 8. Two ingredients for the main construction
In the sequel, $a=2b$ is a positive integer and $V$ is a set of cardinality $a$. We fix a simple graph ${\mathcal{G}}=(V,E)$ belonging to the class ${\mathfrak{B}}_{a}$ defined in Section 1; in particular, $|E|=3a/2$. It is assumed that $V=\{v_{1},\ldots,v_{a}\}$. The vertices of the parts $V_{1}$ and $V_{2}$ of the (bipartite) graph $\mathcal{G}$} are denoted by $u_{1}:=v_{1},\ldots,u_{b}:=v_{b}$ and $w_{1}:=v_{b+1},\ldots,w_{b}:=v_{a}$, respectively.
The graph $\mathcal{G}$} being cubic and bipartite admits a $1$-factorization $F_{\mathcal{G}}$, i.e., the set of three pairwise disjoint matchings the union of which is equal to $E$. It is convenient to index the matchings of $F_{\mathcal{G}}$ by the nonidentity elements of the Klein group $K=C_{2}\times C_{2}$; we set $F_{\mathcal{G}}=\{E_{k}:\ k\in K^{\#}\}$, where $K^{\#}=K\setminus\{1\}$ and $E_{k}$ is the matching indexed by $k$.
8.1. A coherent configuration associated with $\mathcal{G}$}
With each vertex $v\in V$ of the graph $\mathcal{G}$}, we associate a copy $\Omega_{v}$ of the Klein group $K$ and, for a subset $V^{\prime}\subseteq V$, we denote by $\Omega_{V^{\prime}}$ the union of $\Omega_{v}$, $v\in V^{\prime}$. When $V^{\prime}=\{v,v^{\prime}\}$, we abbreviate $\Omega_{v,v^{\prime}}:=\Omega_{\{v,v^{\prime}\}}$.
@@ 8.1. A coherent configuration associated with $\mathcal{G}$}
is a permutation group on the set $\Omega=\Omega_{V}$. It was proved in [8, Section 5] that there exists a coherent configuration ${\mathcal{X}}={\mathcal{X}}({\mathcal{G}},F_{\mathcal{G}})=(\Omega,S)$ that satisfies the following conditions:
@@
    $F({\mathcal{X}})=\{\Omega_{v}:\ v\in V\}$, and also $\operatorname{Aut}({\mathcal{X}})^{\Omega_{v}}=\bar{K}_{v}$ for all $v$,
@@
    the group of all $f\in\operatorname{Iso}({\mathcal{X}})$ such that ${\Omega_{v}}^{f}=\Omega_{v}$ and $f^{\Omega_{v}}\in\bar{K}_{v}$ for all $v\in V$, coincides with $K_{V}$,
@@
    $\operatorname{Aut}({\mathcal{X}})^{\Omega_{v,v^{\prime}}}=\bar{K}_{v}\times\bar{K}_{v^{\prime}}$ for all distinct $v,v^{\prime}\in V$ nonadjacent in $\mathcal{G}$},
@@
    $\operatorname{Aut}({\mathcal{X}})^{\Omega_{v,v^{\prime}}}=\{(x,x^{\prime})\in\bar{K}_{v}\times\bar{K}_{v^{\prime}}:\ \bar{\rho}_{v}(x)=\bar{\rho}_{v^{\prime}}(x^{\prime})\},$
@@
    if $\{u,w\}\in E$ and $\psi\in\operatorname{Sym}(S)$ are such that $\psi(s)\neq s$ if and only if $s\in S_{\Omega_{u},\Omega_{w}}\,\cup\,S_{\Omega_{w},\Omega_{u}}$, then $\psi$ is an algebraic automorphism of $\mathcal{X}$}.
The details concerning the construction and other properties of the coherent configuration $\mathcal{X}$} can also be found in [5, Subsections 4.1.3 and 4.2.1], where $\mathcal{X}$} is defined as a special case of a cubic Klein configuration.
To state the main fact about the coherent configuration $\mathcal{X}$}, that we need later, we fix an edge $e_{0}\in E$ and set $\psi=\psi_{e_{0}}$ to be the algebraic isomorphism in condition (K5) for $e=e_{0}$. Next, for an integer $m\geq 1$, we put
$F_{m}=F_{m}({\mathcal{X}})=\{\Omega_{V^{\prime}}:\ |V^{\prime}|\leq m\}.$
A family $\{f_{\Delta}\in K_{V}:\Delta\in F_{m}\}$ is called an $m$-local system of isomorphisms for $\mathcal{X}$} with respect to $\psi$ if for all $\Delta,\Delta^{\prime}\in F_{m}$, we have
$(f_{\Delta})^{\Delta}\in\operatorname{Iso}({\mathcal{X}}_{\Delta},\psi_{\Delta})\quad\text{and}\quad(f_{\Delta}{f_{\Delta^{\prime}}}^{-1})^{\Delta\cap\Delta^{\prime}}\in\operatorname{Aut}({\mathcal{X}})^{\Delta\cap\Delta^{\prime}}.$
@@
In the above notation, assume that each separator [^5] A separator of the graph $\mathcal{G}$} is a subset $V^{\prime}\subset V$ satisfying the following condition: each connected component of the graph obtained from $\mathcal{G}$} by removing all the vertices of $V^{\prime}$ contains at most half vertices of $\mathcal{G}$}. of the graph $\mathcal{G}$} has cardinality greater than $3m$. Then $\operatorname{Iso}({\mathcal{X}},\psi)=\varnothing$ and there is an $m$-local system of isomorphisms for $\mathcal{X}$} with respect to $\psi$.
@@
8.2. A coset S-ring associated with the graph $\mathcal{G}$}